### v2

# Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities

Abstract
We investigate strong data processing inequalities (SDPIs) of the Rényi-divergence between two discrete distributions when both distributions are passed through a fixed channel. We provide a condition on the channel for which the DPI holds with equality given two arbitrary distributions in the probability simplex. Motivated by this, we examine the contraction behavior for restricted sets of prior distributions via $f$-divergence inequalities: We provide an alternative proof of the optimal reverse Pinsker’s inequality for Rényi-divergences first shown by Binette [1]. We further present an improved Pinsker’s inequality for Rényi-divergence based on the joint range technique by Harremoës and Vajda [2]. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By framing these inequalities in a cross-channel setting, we arrive at SDPIs that can be adapted to use-case specific restrictions of input distribution and channel. We apply these results to the Rényi local differential privacy amplification through post-processing by channels that satisfy no local differential privacy guarantee.
I Introduction
[…]
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\biggl{[}\biggl{(}\frac{dP}{dQ}\biggr{)}^{\alpha}\biggr{]}$ (1)
in an AWGN channel is derived in [10] via the Hockey-stick divergence $E_{\gamma}$. Recently, Jin et al. [11] show that the usual upper bound on contraction via total variation fails to hold for Rényi-divergences. They however prove that the universal $\chi^{2}$ lower bound contraction holds for general channels. Zamanlooy and Asoodeh [6] show that if a channel satisfies local differential privacy (LDP), SDPIs can be derived in terms of a channels LDP guarantee. These bounds are further applied to mixing times in [12]. Related, it is shown in [13] that the LDP guarantee of a channel is closely related to its $f$-divergence contraction properties.
Information-theoretic tools for assessing privacy in data processing systems have highlighted the crucial role of the Rényi-divergence in this context: On one hand, Rényi differential privacy (RDP) [14] has recently been adopted as a powerful relaxation of differential privacy. Since, it has been shown to be highly useful for, e.g, composition and privacy accounting [15, 16]. On the other hand, information-theoretic measures that aim to quantify privacy in terms of inference limits can often be shown to reduce to formulations involving Rényi-divergences. An early example of this is the usage of Rényi-divergence of order infinity in the framework of quantitative information flow [17, 18]. More recently, the highly related notion of maximal leakage [19] and its variants and generalizations [20, 21, 22, 23, 24] emerged, and solidify the important role of Rényi-divergence for adversarial approaches to privacy. Interestingly, the framework suggested in [25] unifies many of the above mentioned approaches to quantifying privacy, including maximal leakage, RDP and, in a local setting, Rényi local differential privacy (RLDP). On a related note, the recently proposed pointwise adaption of maximal leakage, pointwise maximal leakage (PML) [21] is shown in [26] to enable inferential guarantees that determine the privacy risk of outcomes by evaluating the Rényi-divergence of order $\infty$ between the prior and posterior distribution of the secret.
Rényi-divergence is itself not an $f$-divergence. However, it can be shown that the Rényi-divergence of any order $\alpha$ can be written as a monotone transform of an $f$-divergence parameterized by $\alpha$ [27]. Using this result, standard $f$-divergence inequalities can be applied to Rényi-divergence. Numerous such inequalities are derived in [27] and subsequent works. Perhaps closest to the approach used in this work, such inequalities are used to derive SDPIs in [28].
While the contraction of $f$-divergences–and hence of Rényi-divergences–of LDP channels is studied in works like [6, 29, 30, 31], there are fewer results that consider the contraction of channels that satisfy no LDP guarantee. For such channels, the aforementioned PML framework continues to allow for a (finite) privacy quantification. Since both frameworks rely on randomization for privacy protection, it is reasonable to assume that PML exhibits a similar relationship between contraction and privacy guarantee as observed for LDP.
Contributions: In this work, we derive SDPI-style bounds for Rényi-divergences aimed at information-theoretic privacy problems. We present a condition on the channel under which it does not contract the Rényi-divergence in general in Section III. In Section IV, we present tools for bounding Rényi-divergences in terms of total variation distance: We present an improved (and often tight) Pinsker’s inequality for Rényi-divergences given $\alpha>1$, as well as an alternative proof to the optimal reverse Pinsker’s inequality for Rényi-divergence first shown in [1]. We apply these tools to obtain results on the data processing properties of Rényi-divergence for discrete channels in Section V. The results show that the restriction of the input distributions can strongly improve the ability to obtain meaningful bounds on contraction. Finally, we illustrate the results by providing an example use-case that shows how these bounds enable an improved assessment of the RLDP privacy amplification through post-processing. The privacy properties of the non-RLDP channel in this bound are quantified by its PML. This dependence establishes a connection between the contraction of Rényi-divergence and the PML privacy parameter. Further, it shows how privacy assessment with LDP and PML can be used synergistically.
II Preliminaries
We use uppercase letters to denote random variables, lowercase letters to denote their realizations and caligraphic letters to denote sets. Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a transition kernel $P_{Y|X}$ (also referred to as channel or mechanism in the following) that induces a random variable $Y$ at its output. Let $P_{XY}$ denote the joint distribution of $X$ and $Y$. Then we use $P_{XY}=P_{Y|X}\times P_{X}$ to imply that $P_{XY}(x,y)=P_{Y|X=x}(y)P_{X}(x)$ and $P_{Y}=P_{Y|X}\circ P_{X}$ to denote the marginalization $P_{Y}(y)=\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)$. We write $\mathcal{P}_{\mathcal{X}}$ for the probability simplex on the set $\mathcal{X}$. For $N\in\mathbb{N}$, we define $[N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$. Finally, $\log(\cdot)$ denotes the natural logarithm.
II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [32] as
$D_{f}(P||Q)=\mathbb{E}_{Q}\bigg{[}f\bigg{(}\frac{dP}{dQ}\bigg{)}\bigg{]}.$ (2)
The Rényi-divergence [8] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [3]:
Definition 1 ($f_{\alpha}$-divergence).
For any $\alpha>0$, the $f_{\alpha}$-divergence is defined as the $f$-divergence $D_{f_{\alpha}}(P||Q)$, where
$f_{\alpha}(x)=\begin{cases}1-x^{\alpha},\text{ if }\alpha<1,\\ x\log x,\text{ if }\alpha=1,\\ x^{\alpha}-1,\text{ if }\alpha>1.\end{cases}$ (3)
The Rényi-divergence of order $\alpha$ is then obtained by
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\biggl{\{}D_{f_{\alpha}}(P||Q)+1\biggr{\}}.$ (4)
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [10, Section 3.2].
Definition 2 ($E_{\gamma}$-divergence, see, e.g., [27]).
Given some parameter $\gamma>0$, the $E_{\gamma}$-divergence between two discrete distributions $P$ and $Q$ is defined as
$E_{\gamma}(P||Q)=\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-\gamma Q(x)|-\frac{1}{2}|1-\gamma|.$ (5)
For twice differentiable $f$, the relation shown in [33],
$D_{f}(P||Q)=\int_{0}^{\infty}E_{\gamma}(P||Q)f^{\prime\prime}(\gamma)d\gamma,$ (6)
can then be used to obtain any such $D_{f}$-divergence. For $\gamma=1$, the $E_{\gamma}$-divergence is equal to the total variation distance between the two distributions, i.e., $E_{1}(P||Q)=\text{TV}(P||Q)$.
II-B Strong Data Processing Inequalities
All $f$-divergences (including the $f_{\alpha}$-divergence)–and hence also the Rényi-divergence–satisfy the data processing inequality (DPI). The DPI states that for any two probability measures $P_{X},Q_{X}$ on $\mathcal{X}$, and some transition kernel $P_{Y|X}$,
$D_{f}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})\leq D_{f}(P_{X}||Q_{X}).$ (7)
We define the contraction coefficient $\eta_{\alpha}(P_{Y|X})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality.
$\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P})$ (8)
$=$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
$\eta_{\alpha}(P_{Y|X},\mathcal{P})=\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
Here, $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ denotes a subset of the probability simplex on the alphabet $\mathcal{X}$. In this framework, the DPI states that $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$. We are then interested in the conditions for which $\eta_{\alpha}(P_{Y|X},\mathcal{P})<1$, that is, cases in which the inequality in (7) is strict. In this case, we say a strong data processing inequality (SDPI) holds. Similarly to the above, we define $\eta_{\text{TV}}(P_{Y|X},\mathcal{P})$ as the contraction coefficient of the total variation distance given that input distributions are picked from the set $\mathcal{P}$. In the most general case, an SDPI considers $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. However, we will later see that without any restriction, $\eta_{\alpha}(P_{Y|X})=1$ in many cases. We will therefore derive bounds for cases in which $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$. If the uncertainty set $\mathcal{P}$ is clear from context, we omit it for ease of notation.
III Conditions for $\eta_{\alpha}(P_{Y|X})=1$
The following result provides a condition on the channel $P_{Y|X}$ for which there is no contraction, that is, for which no strong data processing inequality holds, in the case $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$.
Proposition 3.
Consider a discrete channel $P_{Y|X}$. Whenever
$\exists x,x^{\prime}\in\mathcal{X}:\textnormal{supp}\big{(}P_{Y|X=x}\big{)}\cap\textnormal{supp}\big{(}P_{Y|X=x^{\prime}}\big{)}=\emptyset,$ (11)
we have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
Proof:
See Appendix A. ∎
III-A Connection to the Confusion Graph of $P_{Y|X}$
Proposition 3 states that a channel does not contract the Rényi-divergence whenever there exist two input symbols that cause conditional distributions with disjoint support at the channel output. This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [34]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
$W=\{(x_{i},x_{j})\in\mathcal{X}^{2}\mid\exists\,y\in\mathcal{Y}:x_{i}\in\mathcal{X}^{(y)}\wedge x_{j}\in\mathcal{X}^{(y)}\},$ (12)
where $\mathcal{X}^{(y)}$ for each $y$ denotes the positions at which the channel matrix contains non-zero entries, that is,
$\mathcal{X}^{(y)}\coloneqq\{x\in\mathcal{X}:P_{Y|X=x}(y)>0\}.$ (13)
Heuristically, an edge $(x_{i},x_{j})$ between input symbols in the confusion graph indicates that there exists an output symbol $y\in\mathcal{Y}$ that can be caused by both $x_{i}$ or $x_{j}$ at the channel input. That is, a missing edge (compared to the complete version) indicates the existence of a pair of symbols that can be sent over the channel without being confused with each other. The confusion graph–along with its graph-theoretic dual–is extensively used in zero-error information theory, see e.g. [35, 36] for a detailed account of results and methods.
Note that a fully connected confusion graph is not necessarily equivalent to a strictly positive channel matrix. Proposition 3 in a loose sense states that for sufficiently sparse $P_{Y|X}$, Rényi-divergences do not contract in general. The following examples illustrates this condition on the channel matrices.
Example 4 (Block-diagonal matrices).
Any block-diagonal matrix $\text{diag}(B_{1},\dots,B_{m})$ with each $B_{i}\in\mathbb{R}^{k_{i}\times k_{i}},\,k_{i}\in\mathbb{N}$, satisfies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
Example 5 ($k$-singular channels).
Inspired by [37], we say a $N\times N$ channel $P_{Y|X}$ is $k$-singular if
$P_{Y|X=x}(y)=\begin{cases}\frac{1}{k},\text{ if }x\in\mathcal{X}^{(y)}\\ 0,\text{ otherwise.}\end{cases}$ (14)
By Example 4, any $k$-singular channel with $k\leq N/2$ satisfies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. For illustration, consider the channels
$P=\begin{bmatrix}\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0\\ \nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}\\ \nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\\ 0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\end{bmatrix},\quad Q=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}.\end{bmatrix}$ (15)
Here, $P$ is $3$-singular and $Q$ is $2$-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
IV Pinsker-type Inequalities for Rényi-divergence
The result in Section III shows that in many cases, the contraction coefficient can not be smaller than one whenever $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. This motivates us to investigate the behavior of $\eta_{\alpha}$ in cases where $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$. As we will see in Section VI, this can be useful in cases in which partial knowledge about the input distributions to a channel is available.
[…]
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [1, Theorem 1]).
Let $\alpha>1$. For any two distributions $P_{X},Q_{X}\in\mathcal{P}$, define
$\Gamma_{\max}\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (16)
$\Gamma_{\min}\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (17)
Then, the maximum $f_{\alpha}$-divergence corresponding to the $\alpha$-Rényi-divergence $D_{\alpha}(P_{X}||Q_{X})$ between the two distributions is upper bounded by
$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\text{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max},\Gamma_{\min}),$ (18)
where
$R_{\alpha}(\Gamma_{\max},\Gamma_{\min})=\frac{\Gamma_{\max}^{\alpha}-1}{\Gamma_{\max}-1}-\frac{1-\Gamma_{\min}^{\alpha}}{1-\Gamma_{\min}}.$ (19)
We provide an alternative proof to [1]. Our proof utilizes the $E_{\gamma}$ identity of $f$-divergences.
Proof:
Fix $Q_{X}$ and $P_{X}$ and let $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$. The proof depends on the fact that $E_{\gamma}(P_{X}||Q_{X})$ can only be non-zero on the interval $(\Gamma_{\min},\Gamma_{\max})$. To see this, observe that
$E_{\gamma}(P_{X}||Q_{X})=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|$ (20)
and whenever $P_{X}(x)-\gamma Q_{X}(x)\leq 0$ for all $x\in\mathcal{X}$ or $P_{X}(x)-\gamma Q_{X}(x)\geq 0$ for all $x\in\mathcal{X}$, we have
$E_{\gamma}(P_{X}||Q_{X})$ $=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|$ (21)
$=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{(}\pm P_{X}(x)\mp\gamma Q_{X}(x)\biggr{)}-\frac{1}{2}|1-\gamma|$ (22)
$=\frac{1}{2}|1-\gamma|-\frac{1}{2}|1-\gamma|=0.$ (23)
This happens exactly when $\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(P_{X},Q_{X})$ or $\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\coloneqq\Gamma_{\max}(P_{X},Q_{X})$. Hence, $E_{\gamma}$ is zero everywhere but on the interval $\gamma\in(\Gamma_{\min}(P_{X},Q_{X}),\Gamma_{\max}(P_{X},Q_{X}))\eqqcolon I_{\geq 0}$. From this, we get
$E_{\gamma}(P_{X}||Q_{X})=\begin{cases}0,&\text{ if }\gamma\notin I_{\geq 0},\\ u(\gamma),&\text{ if }\gamma\in I_{\geq 0}.\end{cases}$ (24)
What remains is to bound $u(\gamma)$ to obtain an upper bound in (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [10]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $u(\gamma)$ by the piecewise linear function
$u(\gamma)\leq g(\gamma)\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\ \Delta\frac{\gamma-\Gamma_{\min}(P_{X},Q_{X})}{1-\Gamma_{\min}(P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\ \Delta\frac{-\gamma+\Gamma_{\max}(P_{X},Q_{X})}{\Gamma_{\max}(P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (25)
The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$ and solving
[…]
which yields a bound on the $f_{\alpha}$-divergence by (6). ∎
Note that this proof technique can also be used to obtain the more general result in [1] for any $f$-divergence for which $f$ is twice differentiable by replacing $\gamma^{\alpha-2}$ with the general $f^{\prime\prime}(\gamma)$ in (26). Further, if more information about the distributions is known, more specific bounds on $f(\gamma)$ can potentially be used to obtain more specific inequalities in these cases.
Remark 7.
The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}$. To see this, let again $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$ and observe that we can plug in the result form Theorem 6 into (4) to obtain
$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})+1\big{)}.$ (27)
Since $\Gamma_{\min}\leq 1$ and $\Gamma_{\max}\geq 1$, we get the limit as
$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})\biggr{)}$ (28)
$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}=\log\Gamma_{\max}.$ (29)
Further, the bound in [27, Theorem 35] can be obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}=0$.
Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergences of orders $\alpha>1$.
Theorem 8 (Pinsker’s inequality for Rényi-divergences).
For two mutually absolutely continuous probability distributions $P_{X}$ and $Q_{X}$ and any $\alpha>1$, we have
$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl{(}\text{TV}(P_{X}||Q_{X})\biggr{)}$ (30)
where
$g_{\alpha}(t)=\begin{cases}e^{2(\alpha-1)t^{2}}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha<2,\\ (4t^{2}+1)^{\alpha-1}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha\geq 2,\\ (1-t)^{1-\alpha}-1,&\text{ otherwise }\end{cases}$ (31)
Proof:
The proof follows from bounding the joint range [2] of the two divergences and is presented in Appendix B. ∎
Remark 9.
For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [3]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$.
Remark 10.
The function $g_{\alpha}$ is injective, but not bijective. That is, in order to obtain an inverse bound (later used in Corollary 11) we can define the piecewise function
$g^{-1}_{\alpha}(s)=\begin{cases}\sqrt{\frac{\log(s+1)}{2(\alpha-1)}},&\text{if}\alpha<2,s<h_{1}(\alpha),\\ \frac{1}{2}\sqrt{(s+1)^{(\frac{1}{\alpha-1})}-1},&\text{if }\alpha\geq 2,s<h_{2}(\alpha)\\ \max\{1-(s+1)^{\frac{1}{1-\alpha}},\nicefrac{{1}}{{\alpha}}\},&\text{otherwise},\end{cases}$ (32)
where $h_{2}(\alpha)\coloneqq(1+4\alpha^{-2})^{\alpha-1}-1$ and $h_{1}(\alpha)\coloneqq 2-\nicefrac{{2}}{{\alpha}}$.
An example of the the presented bound on the joint range for $\alpha=4$ is shown in Figure 1. Other values of $\alpha$ yield a visually similar result, with the discontinuity at $\nicefrac{{1}}{{\alpha}}$ moving closer to the origin as the value of $\alpha$ increases.
Figure 1: Joint range of $D_{f_{\alpha}}(P||Q)$ and $\text{TV}(P||Q)$ for the case $\alpha=4$ as well as the bound presented in Theorem 8.
V Strong Data Processing with Pinsker
[…]
$\Gamma_{\max}=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (33)
$\Gamma_{\min}=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (34)
[…]
Corollary 11.
For a discrete channel $P_{Y|X}$, a set $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ and some $\alpha>1$, let $\eta\coloneqq\eta_{TV}(P_{Y|X})$. We have
$D_{f_{\alpha}}(P_{Y}||Q_{Y})\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(D_{f_{\alpha}}(P_{X}||Q_{X})).$ (35)
Proof:
Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$ and the inequality in Theorem 8. ∎
This result stresses the relation between $\Gamma_{\max}$ and $\Gamma_{\min}$ as formulated in (33) and (34) and the contraction behavior of a channel: If $\Gamma_{\max}-\Gamma_{\min}$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}$ and $\Gamma_{\min}$ such that an operational bound is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. We present one such use-case specific bound in the following.
VI Application: Post-Processing for RLDP
In this section, we show how the above results can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [14, 25]. Consider the Markov chain $W-X-Y$, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, i.e.,
$\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (36)
Further, assume that the transition kernel $P_{Y|X}$ does not satisfy any finite RLDP guarantee, that is, the channel matrix contains zero elements. From the post-processing property of RLDP, it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there exist channels $P_{Y|X}$ containing zero elements that will result in an improved guarantee of the composed system $P_{Y|W}$ if the output of the RLDP $P_{X|W}$ is applied to their input. In fact, using the results derived above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
$D_{f_{\alpha}}(P_{Y|W=w}||P_{Y|W=w^{\prime}})$ (37)
$=D_{f_{\alpha}}(P_{Y|X}\circ P_{X|W=w}||P_{Y|X}\circ P_{X|W=w^{\prime}})$ (38)
$\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}\biggl{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\biggr{)},$ (39)
where $\eta\coloneqq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})$ is again a shorthand for the contraction coefficient of the total variation distance. Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can therefore bound the RLDP guarantee of the entire system by applying (4) as
$\varepsilon^{*}\coloneqq\text{RLDP}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (40)
where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence,
$\varphi(\varepsilon_{f},\alpha)\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (41)
Further, in this setup, the set of input distributions $\mathcal{P}$ is just the set of rows of the mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$. Hence, the quantities $\Gamma_{\max}$ and $\Gamma_{\min}$ reduce to
$\Gamma_{\max}=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (42)
$\Gamma_{\min}=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (43)
While these quantities can be computed directly if both channels are perfectly known, it is useful to obtain a bound that only depends on the privacy guarantees of the channels. The following proposition provides such a bound, where the privacy of the second, non-RLDP channel is quantified by its pointwise maximal leakage (PML) [21]. The proposition shows one way in which the increased flexibility of leakage assessment with PML can complement (L)DP guarantees.
Proposition 12.
If $P_{X|W}$ is a $N\times N$ randomized response mechanisms satisfying $\varepsilon$-LDP according to [31], then
$\Gamma_{\max}\leq\biggl{(}1-\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (44)
$\Gamma_{\min}\geq\biggl{(}1+\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (45)
where $\ell_{w}(X\to y)$ denotes the PML from $X$ to an outcome $Y=y$ under the input distribution induced by $W=w$, i.e.,
$\ell_{w}(X\to y)\coloneqq\log\frac{\max_{x\in\mathcal{X}}P_{Y|X=x}(y)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}.$ (46)
Further, equality is achieved in both bounds if $P_{Y|X}$ has a zero entry in each column, i.e., if $\min_{x}P_{Y|X=x}(y)=0$ $\forall y\in\mathcal{Y}$.
Proof:
See Appendix C. ∎
The following example illustrates this bound.
Figure 2: RLDP and corresponding bound in Example 13.
Example 13.
Let $P_{X|W}$ be a $5\times 5$ randomized response mechanism in [31] satisfying $\log 6$-LDP, and $P_{Y|X}$ given by
$P_{Y|X}=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0&0\\ 0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0\\ 0&0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ \nicefrac{{1}}{{2}}&0&0&0&\nicefrac{{1}}{{2}}\end{bmatrix}.$ (47)
[…]
Appendix A Proof of Proposition 3
For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have
$\eta_{\alpha}(P_{Y|X})=$ (48)
$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (49)
and the statement follows by constructing $P_{X}$ and $Q_{X}$ with full support such that in the limiting case to the support set boundaries, we get
$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (50)
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, constructing this distribution proves $\eta_{\alpha}(P_{Y|X})=1$.
Construction of $P_{X}$ and $Q_{X}$: For ease of notation, let $\mathcal{X}\coloneqq[N]$. Further, define the indicator function by
$\delta_{i}(x)=\begin{cases}1,&\text{if }x=i\\ 0,&\text{otherwise}.\end{cases}$ (51)
For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
$P_{X}^{\gamma}(x)=(1-\gamma)\delta_{i}(x)+\frac{\gamma}{N-1}(1-\delta_{i}(x)),$ (52)
and let
$Q_{X}^{\gamma}(x)=\frac{1-\gamma}{2}(\delta_{i}(x)+\delta_{j}(x))+\frac{\gamma}{N-2}(1-\delta_{i}(x)-\delta_{j}(x)),$ (53)
where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Then we get
$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ $=\frac{\log\biggl{\{}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (54)
$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (55)
$=\frac{(1-\alpha)\log Q^{0}_{X}(i)}{(1-\alpha)\log Q^{0}_{X}(i)}=1,$ (56)
where the second to last equality follows from the row-stochasticity of $P_{Y|X}$.
By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$. Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$. Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof. ∎
Appendix B Proof of Theorem 8
From [27] and the fact that $D_{f_{\alpha}}$ is a monotone transform of the Rényi-divergence of the same order, we know that
$\inf_{P,Q\in\mathcal{P}_{X}:\,TV(P||Q)=t}D_{f_{\alpha}}(P||Q)=\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q),$ (57)
where $d_{f_{\alpha}}(p||q)$ denotes the binary $f_{\alpha}$-divergence
$d_{f_{\alpha}}(p||q)\coloneqq\frac{p^{\alpha}}{q^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-q)^{\alpha-1}}-1.$ (58)
From the definition of the total variation, we get
$\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\biggl{\{}\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}\biggr{\}}-1.$ (59)
Define
$\psi_{\alpha}(p,t)\coloneqq\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}-1.$ (60)
Taking derivatives with respect to $p$ yields
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}$ (61)
and for any $\alpha>1$ and all $t<1,\,p\leq 1-t$,
$\frac{\partial^{2}}{\partial p^{2}}\psi_{\alpha}(p,t)=(\alpha-1)\alpha t^{2}\biggl{[}\frac{p^{\alpha-2}}{(p+t)^{\alpha+1}}+\frac{(1-p)^{\alpha-2}}{(1-p-t)^{\alpha+1}}\biggr{]}\geq 0,$ (62)
from which we can conclude that $\psi_{\alpha}(p,t)$ is convex in $p$.
First, consider the case $t\geq\nicefrac{{1}}{{\alpha}}$. We can bound the first derivative by
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)\geq(\alpha t+p-1)\biggl{(}\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{)}\geq 0,$ (63)
with equality for $t=\nicefrac{{1}}{{\alpha}}$ if $p=0$. Otherwise strict inequality holds. Together with the convexity of $\psi_{\alpha}$ in $p$, this implies that the maximum value of $\psi_{\alpha}$ for all $t\geq\nicefrac{{1}}{{\alpha}}$ will be achieved at $p=0$. At this point, we have
$\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}.$ (64)
Secondly, for $t<\nicefrac{{1}}{{\alpha}}$, we first consider $\alpha\geq 2$. Let $P,Q$ be binary probability distributions with $P(0)=p$ and $Q(0)=p+t$ Then, noticing that $x\mapsto x^{\alpha-1}$ is convex for $\alpha\geq 2$, we apply Jensen’s inequality to obtain
$\psi_{\alpha}(p,t)$ $=\mathbb{E}_{X\sim P}\biggl{[}\biggl{(}\frac{P(X)}{Q(X)}\biggr{)}^{\alpha-1}-1\biggr{]}$
$\geq\biggl{(}\mathbb{E}_{X\sim P}\biggl{[}\frac{P(X)}{Q(X)}\biggr{]}\biggr{)}^{\alpha-1}-1$
$=\biggl{(}1+\frac{t^{2}}{(p+t)(1-p-t)}\biggr{)}^{\alpha-1}-1$
$\geq(1+4t^{2})^{\alpha-1}-1.$
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [9], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$)
$\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$ (65)
will continue to hold for $\alpha>1$. Utilizing (4), this yields
$D_{f_{\alpha}}(P||Q)\geq\exp\big{(}2(\alpha-1)\text{TV}(P||Q)^{2}\big{)}-1.$ (66)
Putting the derived bounds together to form $g_{\alpha}(t)$ finishes the proof. ∎
Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}=0.$ (67)
In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$. Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by
$g_{\alpha}(t)=\psi_{\alpha}(p^{*},t).$ (68)
However, it remains unclear if $p^{*}$ can be expressed in closed form. Due to this complication, the result above utilizes more rudimentary bounds in this case.
Appendix C Proof of Proposition 12
Note that for a randomized response mechanism according to [31], we can write for any two rows $w$, $w^{\prime}$,
$P_{X|W=w}(x)=P_{X|W=w^{\prime}}(x)+\kappa(x),$ (69)
where
$\kappa(x)\in\{0,\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1},\frac{1-e^{\varepsilon}}{|\mathcal{X}|+e^{\varepsilon}-1}\}$ (70)
and each non-zero element is taken exactly once. Therefore,
$||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}.$ (71)
Using this observation, we can write
$\Gamma_{\max}$ $=\max_{w,w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w}(x)}$ (72)
$=\max_{\kappa:\,||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}}\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)+\sum_{x}P_{Y|X=x}(y)\kappa(x)}$ (73)
$=\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\biggl{(}\max_{x}P_{Y|X=x}(y)-\min_{x}P_{Y|X=x}(y)\biggr{)}}$ (74)
$\leq\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\max_{x}P_{Y|X=x}(y)}.$ (75)
Now, fix $y$ and let $w^{*}\in\mathcal{W}$ be the solution to the maximization over $\mathcal{W}$. By letting
$\ell_{w^{*}}(X\to y)=\log\frac{\max_{x}P_{Y|X=x}(y)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{*}}(x)},$ (76)
we can further simplify the above bound as
$\Gamma_{\max}\leq\max_{y}\frac{1}{1-\frac{||\kappa||_{1}}{2}\exp\ell_{w^{*}}(X\to y)}=\biggl{(}1-\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1}.$ (77)
The proof for $\Gamma_{\min}$ follows analogously. ∎

### v3

@@
-Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities
+Bounds on the Privacy Amplification of Arbitrary Channels via the Contraction of $f_{\alpha}$-Divergence
 Abstract
-We investigate strong data processing inequalities (SDPIs) of the Rényi-divergence between two discrete distributions when both distributions are passed through a fixed channel.
-We provide a condition on the channel for which the DPI holds with equality given two arbitrary distributions in the probability simplex.
-Motivated by this, we examine the contraction behavior for restricted sets of prior distributions via $f$-divergence inequalities: We provide an alternative proof of the optimal reverse Pinsker’s inequality for Rényi-divergences first shown by Binette [1].
-We further present an improved Pinsker’s inequality for Rényi-divergence based on the joint range technique by Harremoës and Vajda [2].
+We examine the privacy amplification of channels that do not necessarily satisfy any LDP guarantee by analyzing their contraction behavior in terms of $f_{\alpha}$-divergence, an $f$-divergence related to Rényi-divergence via a monotonic transformation.
+We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by Harremoës and Vajda [1].
 The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$.
-By framing these inequalities in a cross-channel setting, we arrive at SDPIs that can be adapted to use-case specific restrictions of input distribution and channel.
-We apply these results to the Rényi local differential privacy amplification through post-processing by channels that satisfy no local differential privacy guarantee.
+By applying these inequalities in a cross-channel setting, we arrive at strong data processing inequalities for $f_{\alpha}$-divergence that can be adapted to use-case specific restrictions of input distributions and channel.
+The application of these results to privacy amplification shows that even very sparse channels can lead to significant privacy amplification when used as a post-processing step after local differentially private mechanisms.
 I Introduction
-[…]
-$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\biggl{[}\biggl{(}\frac{dP}{dQ}\biggr{)}^{\alpha}\biggr{]}$ (1)
-in an AWGN channel is derived in [10] via the Hockey-stick divergence $E_{\gamma}$.
+An important property of local differential privacy (LDP) is its invariance to post-processing.
+Informally, the post-processing property states that whatever further processing is applied to the output of a system satisfying LDP will not weaken the overall system’s privacy guarantee.
+Mathematically, this property can be seen as a form of data processing inequality (DPI).
+As an extension of this observation, it is natural to ask about the privacy amplification of a post-processing operation.
+Basic results state that if a post-processing operation satisfying $\varepsilon_{2}$-LDP is applied to a system satisfying $\varepsilon_{1}$-LDP, the composition of both systems will satisfy at least $\min\{\varepsilon_{1},\varepsilon_{2}\}$-LDP.
+However, it is generally not know when arbitrary post-processing operations, that is, operations that themselves do not satisfy any LDP guarantee, may improve the overall systems LDP guarantee.
+We take steps to answer this question in this paper.
+In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [2, Chapter 33]).
+For pure LDP, the divergence measure of interest is the Rényi-divergence [3, 4] of order $\infty$, since
+$\max_{y}\log\frac{P_{X|W=w}(y)}{P_{X|W=w^{\prime}}(y)}=\lim_{\alpha\to\infty}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}}).$ (1)
+Gilani et al.
+[5] suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$.
+With this framework at hand, studying the privacy amplification of a channel becomes equivalent to studying SDPIs for Rényi-divergences.
+While SDPIs have been extensively studied for $f$-divergences, results on Rényi-divergences are so far less explored.
+Although Rényi-divergence is not an $f$-divergence, it can be expressed as a monotonic transformation of a corresponding $f_{\alpha}$-divergence [6], a family of $f$-divergences parameterized by $\alpha$.
+This connection allows us to translate results on $f$-divergences to Rényi-divergences—results that are often easier to obtain.
+I-A Contributions
+In this work, we derive several technical tools for studying the privacy amplification of arbitrary discrete channels—those not necessarily satisfying any LDP guarantee—by analyzing how they affect the Rényi-LDP of composed systems.
+Our key contributions are:
+  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on privacy amplification in Section III.
+  • We develop new tools for bounding the $f_{\alpha}$-divergence in terms of total variation distance in Section IV. Specifically, we provide:
+    • an improved and often tight version of Pinsker’s inequality for $f_{\alpha}$-divergence (for $\alpha>1$);
+    • a simplification of the recently presented proof of Binette’s inequality [7] in [8] based on $E_{\gamma}$-divergence.
+  • Using these tools, we analyze the contraction of $f_{\alpha}$-divergence and show how input distribution constraints can yield meaningful bounds in Section V.
+  • Finally, in Section VI, we apply these bounds to the Rényi-LDP amplification of arbitrary post-processing channels.
+Our results show that many channels—despite not satisfying any (finite) LDP guarantee—can result in significant privacy amplification if applied to the output of a mechanism satisfying LDP.
+We illustrate these findings by applying the proposed bounds to randomized response mechanisms post-processed by channels with sparse channel matrices (i.e., matrices with many zeros).
+I-B Related Work
+The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [2, 9, 10, 11, 12, 13].
+For Rényi-divergences, Polyanskiy and Wu [14] provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$.
 Recently, Jin et al.
-[11] show that the usual upper bound on contraction via total variation fails to hold for Rényi-divergences.
-They however prove that the universal $\chi^{2}$ lower bound contraction holds for general channels.
-Zamanlooy and Asoodeh [6] show that if a channel satisfies local differential privacy (LDP), SDPIs can be derived in terms of a channels LDP guarantee.
+[15] show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid.
+In the context of privacy, Zamanlooy and Asoodeh [11] show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee.
 These bounds are further applied to mixing times in [16].
-Related, it is shown in [13] that the LDP guarantee of a channel is closely related to its $f$-divergence contraction properties.
-Information-theoretic tools for assessing privacy in data processing systems have highlighted the crucial role of the Rényi-divergence in this context: On one hand, Rényi differential privacy (RDP) [14] has recently been adopted as a powerful relaxation of differential privacy.
-Since, it has been shown to be highly useful for, e.g, composition and privacy accounting [15, 16].
-On the other hand, information-theoretic measures that aim to quantify privacy in terms of inference limits can often be shown to reduce to formulations involving Rényi-divergences.
-An early example of this is the usage of Rényi-divergence of order infinity in the framework of quantitative information flow [17, 18].
-More recently, the highly related notion of maximal leakage [19] and its variants and generalizations [20, 21, 22, 23, 24] emerged, and solidify the important role of Rényi-divergence for adversarial approaches to privacy.
-Interestingly, the framework suggested in [25] unifies many of the above mentioned approaches to quantifying privacy, including maximal leakage, RDP and, in a local setting, Rényi local differential privacy (RLDP).
-On a related note, the recently proposed pointwise adaption of maximal leakage, pointwise maximal leakage (PML) [21] is shown in [26] to enable inferential guarantees that determine the privacy risk of outcomes by evaluating the Rényi-divergence of order $\infty$ between the prior and posterior distribution of the secret.
-Rényi-divergence is itself not an $f$-divergence.
-However, it can be shown that the Rényi-divergence of any order $\alpha$ can be written as a monotone transform of an $f$-divergence parameterized by $\alpha$ [27].
-Using this result, standard $f$-divergence inequalities can be applied to Rényi-divergence.
-Numerous such inequalities are derived in [27] and subsequent works.
-Perhaps closest to the approach used in this work, such inequalities are used to derive SDPIs in [28].
-While the contraction of $f$-divergences–and hence of Rényi-divergences–of LDP channels is studied in works like [6, 29, 30, 31], there are fewer results that consider the contraction of channels that satisfy no LDP guarantee.
-For such channels, the aforementioned PML framework continues to allow for a (finite) privacy quantification.
-Since both frameworks rely on randomization for privacy protection, it is reasonable to assume that PML exhibits a similar relationship between contraction and privacy guarantee as observed for LDP.
-Contributions: In this work, we derive SDPI-style bounds for Rényi-divergences aimed at information-theoretic privacy problems.
-We present a condition on the channel under which it does not contract the Rényi-divergence in general in Section III.
-In Section IV, we present tools for bounding Rényi-divergences in terms of total variation distance: We present an improved (and often tight) Pinsker’s inequality for Rényi-divergences given $\alpha>1$, as well as an alternative proof to the optimal reverse Pinsker’s inequality for Rényi-divergence first shown in [1].
-We apply these tools to obtain results on the data processing properties of Rényi-divergence for discrete channels in Section V.
-The results show that the restriction of the input distributions can strongly improve the ability to obtain meaningful bounds on contraction.
-Finally, we illustrate the results by providing an example use-case that shows how these bounds enable an improved assessment of the RLDP privacy amplification through post-processing.
-The privacy properties of the non-RLDP channel in this bound are quantified by its PML.
-This dependence establishes a connection between the contraction of Rényi-divergence and the PML privacy parameter.
-Further, it shows how privacy assessment with LDP and PML can be used synergistically.
+Related, it is shown in [17] that the LDP guarantee of a mechanism is closely related to its $f$-divergence contraction properties.
+The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [6] and subsequent works, including an applications to SDPIs in [18].
 II Preliminaries
 We use uppercase letters to denote random variables, lowercase letters to denote their realizations and caligraphic letters to denote sets.
-Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a transition kernel $P_{Y|X}$ (also referred to as channel or mechanism in the following) that induces a random variable $Y$ at its output.
+Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a channel $P_{Y|X}$ (also referred to as mechanism if it satisfies an RLDP guarnatee in the following) that induces a random variable $Y$ at its output.
 Let $P_{XY}$ denote the joint distribution of $X$ and $Y$.
@@
 We write $\mathcal{P}_{\mathcal{X}}$ for the probability simplex on the set $\mathcal{X}$.
-For $N\in\mathbb{N}$, we define $[N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$.
+For $N\in\mathbb{N}$, we define $N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$.
 Finally, $\log(\cdot)$ denotes the natural logarithm.
@@
 $D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\biggl{\{}D_{f_{\alpha}}(P||Q)+1\biggr{\}}.$ (4)
-We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [10, Section 3.2].
+We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [14, Section 3.2].
 Definition 2 ($E_{\gamma}$-divergence, see, e.g., [6]).
@@
 $D_{f}(P||Q)=\int_{0}^{\infty}E_{\gamma}(P||Q)f^{\prime\prime}(\gamma)d\gamma,$ (6)
-can then be used to obtain any such $D_{f}$-divergence.
+can then be used to obtain any such $f$-divergence.
 For $\gamma=1$, the $E_{\gamma}$-divergence is equal to the total variation distance between the two distributions, i.e., $E_{1}(P||Q)=\text{TV}(P||Q)$.
 II-B Strong Data Processing Inequalities
-All $f$-divergences (including the $f_{\alpha}$-divergence)–and hence also the Rényi-divergence–satisfy the data processing inequality (DPI).
+All $f$-divergences (including the $f_{\alpha}$-divergence)—and hence also the Rényi-divergence—satisfy the data processing inequality (DPI).
 The DPI states that for any two probability measures $P_{X},Q_{X}$ on $\mathcal{X}$, and some transition kernel $P_{Y|X}$,
 $D_{f}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})\leq D_{f}(P_{X}||Q_{X}).$ (7)
-We define the contraction coefficient $\eta_{\alpha}(P_{Y|X})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality.
+We define the contraction coefficient $\eta_{\alpha}(P_{Y|X},\mathcal{P})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality, that is,
 $\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P})$ (8)
-$=$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
-$\eta_{\alpha}(P_{Y|X},\mathcal{P})=\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
+$\coloneqq$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
+$\eta_{\alpha}(P_{Y|X},\mathcal{P})\coloneqq\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
 Here, $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ denotes a subset of the probability simplex on the alphabet $\mathcal{X}$.
@@
 In the most general case, an SDPI considers $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$.
-However, we will later see that without any restriction, $\eta_{\alpha}(P_{Y|X})=1$ in many cases.
+However, we will see below that without any restriction, $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$ in many cases.
 We will therefore derive bounds for cases in which $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$.
-If the uncertainty set $\mathcal{P}$ is clear from context, we omit it for ease of notation.
-III Conditions for $\eta_{\alpha}(P_{Y|X})=1$
+III Conditions for $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$
 The following result provides a condition on the channel $P_{Y|X}$ for which there is no contraction, that is, for which no strong data processing inequality holds, in the case $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$.
@@
 Proof:
-See Appendix A. ∎
+We provide a constructive proof in Appendix B.
+Here, we show Proposition 3 by a simple chain of arguments following from previous results.
+Let $(X,Y)\sim P_{Y|X}\times P_{X}$ be arbitrary but fixed.
+Define the maximal correlation as
+$\rho_{m}(X;Y)\coloneqq\sup_{f,g}\mathbb{E}_{(X,Y)\sim P_{XY}}\Big{[}f(X)g(Y)\Big{]},$ (12)
+where the supremum is over all $f:\mathcal{X}\to\mathbb{R}$ and $g:\mathcal{Y}\to\mathbb{R}$ such that
+$\mathbb{E}\big{[}f(X)\big{]}=\mathbb{E}\big{[}g(Y)\big{]}=0,\quad\mathbb{E}\big{[}f^{2}(X)\big{]}=\mathbb{E}\big{[}g^{2}(Y)\big{]}=1.$ (13)
+The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, in particular, it is shown in, e.g., [2, Theorem 33.6(c)] that
+$\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\rho_{m}^{2}(X;Y)\eqqcolon s(P_{Y|X}).$ (14)
+It was shown by Ahlswede and Gács [13] that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$.
+Together with the data processing inequality for $f$-divergences, this implies that whenever (11) holds, we have $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
+Next, it was shown in [15, Corollary 2] that
+$\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\geq\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}}).$ (15)
+By the data processing inequality for Rényi-divergences, we also have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$.
+Hence, if (11) holds, we have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
+∎
 III-A Connection to the Confusion Graph of $P_{Y|X}$
@@
 $P=\begin{bmatrix}\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0\\ \nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}\\ \nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\\ 0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\end{bmatrix},\quad Q=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}.\end{bmatrix}$ (19)
-Here, $P$ is $3$-singular and $Q$ is $2$-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
-IV Pinsker-type Inequalities for Rényi-divergence
+Here, $P$ is 33-singular and $Q$ is 22-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
+IV Pinsker-type Inequalities for $f_{\alpha}$-divergence
 The result in Section III shows that in many cases, the contraction coefficient can not be smaller than one whenever $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$.
@@
 The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions.
-Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [1, Theorem 1]).
+It was first shown by Binette [7] for general $f$-divergences.
+Here, we present a proof that utilizes the integral representation of $f$-divergences via the $E_{\gamma}$-divergence in (6).
+We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche and Tomamichel [8, Proposition 5.2].
+However, the proof presented below slightly simplifies arguments, and we include it here as we believe it to be instructive.
+Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [7, Theorem 1]).
 Let $\alpha>1$. For any two distributions $P_{X},Q_{X}\in\mathcal{P}$, define
-$\Gamma_{\max}\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (16)
-$\Gamma_{\min}\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (17)
-Then, the maximum $f_{\alpha}$-divergence corresponding to the $\alpha$-Rényi-divergence $D_{\alpha}(P_{X}||Q_{X})$ between the two distributions is upper bounded by
-$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\text{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max},\Gamma_{\min}),$ (18)
-where
-$R_{\alpha}(\Gamma_{\max},\Gamma_{\min})=\frac{\Gamma_{\max}^{\alpha}-1}{\Gamma_{\max}-1}-\frac{1-\Gamma_{\min}^{\alpha}}{1-\Gamma_{\min}}.$ (19)
-We provide an alternative proof to [1]. Our proof utilizes the $E_{\gamma}$ identity of $f$-divergences.
+$\Gamma_{\max}(X)\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (20)
+$\Gamma_{\min}(X)\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (21)
+Then, the maximum $f_{\alpha}$-divergence between the two distributions is upper bounded by
+$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\textnormal{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max}(X),\Gamma_{\min}(X)),$ (22)
+where for $u\in[1,\infty)\cup\{\infty\}$ and $v\in[0,1]$,
+$R_{\alpha}(u,v)=\frac{u^{\alpha}-1}{u-1}-\frac{1-v^{\alpha}}{1-v}.$ (23)
 Proof:
 Fix $Q_{X}$ and $P_{X}$ and let $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$.
-The proof depends on the fact that $E_{\gamma}(P_{X}||Q_{X})$ can only be non-zero on the interval $(\Gamma_{\min},\Gamma_{\max})$.
+The proof depends on the fact that the only interval on which we may have $E_{\gamma}(P_{X}||Q_{X})\neq 0$ is $(\Gamma_{\min}(X),\Gamma_{\max}(X))$.
 To see this, observe that
-$E_{\gamma}(P_{X}||Q_{X})=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|$ (20)
+$E_{\gamma}(P_{X}||Q_{X})=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|,$ (24)
 and whenever $P_{X}(x)-\gamma Q_{X}(x)\leq 0$ for all $x\in\mathcal{X}$ or $P_{X}(x)-\gamma Q_{X}(x)\geq 0$ for all $x\in\mathcal{X}$, we have
@@
 $=\frac{1}{2}|1-\gamma|-\frac{1}{2}|1-\gamma|=0.$ (27)
-This happens exactly when $\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(P_{X},Q_{X})$ or $\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\coloneqq\Gamma_{\max}(P_{X},Q_{X})$.
-Hence, $E_{\gamma}$ is zero everywhere but on the interval $\gamma\in(\Gamma_{\min}(P_{X},Q_{X}),\Gamma_{\max}(P_{X},Q_{X}))\eqqcolon I_{\geq 0}$.
+This happens exactly when
+$\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(X,P_{X},Q_{X}),$ (28)
+or,
+$\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\max}(X,P_{X},Q_{X}).$ (29)
+Hence, $E_{\gamma}$ must be zero everywhere but on the interval between these two quantities, that is, the only values of $\gamma$ that allow a non-zero value of $E_{\gamma}$ are
+$\gamma\in(\Gamma_{\min}(X,P_{X},Q_{X}),\Gamma_{\max}(X,P_{X},Q_{X}))\eqqcolon I_{\geq 0}.$ (30)
 From this, we get
-$E_{\gamma}(P_{X}||Q_{X})=\begin{cases}0,&\text{ if }\gamma\notin I_{\geq 0},\\ u(\gamma),&\text{ if }\gamma\in I_{\geq 0}.\end{cases}$ (24)
-What remains is to bound $u(\gamma)$ to obtain an upper bound in (6).
+[…]
+What remains is to bound $E_{\gamma}$ on the interval $I_{\geq 0}$, which yields an upper bound on $D_{f_{\alpha}}$ by (6).
 To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [14].
 Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$.
-By convexity, we can then upper bound $u(\gamma)$ by the piecewise linear function
-$u(\gamma)\leq g(\gamma)\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\ \Delta\frac{\gamma-\Gamma_{\min}(P_{X},Q_{X})}{1-\Gamma_{\min}(P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\ \Delta\frac{-\gamma+\Gamma_{\max}(P_{X},Q_{X})}{\Gamma_{\max}(P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (25)
-The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$ and solving
-[…]
-which yields a bound on the $f_{\alpha}$-divergence by (6). ∎
+By convexity, we can then upper bound $E_{\gamma}$ by the piecewise linear function
+$E_{\gamma}$ $(P_{X}||Q_{X})\leq g(\gamma)$ (32)
+$\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\[6.99997pt] \Delta\frac{\gamma-\Gamma_{\min}(X,P_{X},Q_{X})}{1-\Gamma_{\min}(X,P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\[6.99997pt] \Delta\frac{-\gamma+\Gamma_{\max}(X,P_{X},Q_{X})}{\Gamma_{\max}(X,P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (33)
+The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$, which together with (6) yields
+$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\alpha(\alpha-1)\int_{\Gamma_{\min}(X)}^{\Gamma_{\max}(X)}g(\gamma)\gamma^{\alpha-2}d\gamma,$ (34)
+and the desired statement follows by solving this integral. ∎
 Note that this proof technique can also be used to obtain the more general result in [7] for any $f$-divergence for which $f$ is twice differentiable by replacing $\gamma^{\alpha-2}$ with the general $f^{\prime\prime}(\gamma)$ in (34).
@@
 Remark 7.
-The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}$.
+The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}(X)$.
 To see this, let again $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$ and observe that we can plug in the result form Theorem 6 into (4) to obtain
-$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})+1\big{)}.$ (27)
-Since $\Gamma_{\min}\leq 1$ and $\Gamma_{\max}\geq 1$, we get the limit as
-$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})\biggr{)}$ (28)
-$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}=\log\Gamma_{\max}.$ (29)
-Further, the bound in [27, Theorem 35] can be obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}=0$.
-Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergences of orders $\alpha>1$.
-Theorem 8 (Pinsker’s inequality for Rényi-divergences).
+$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max}(X),\Gamma_{\min}(X))+1\big{)}.$ (35)
+Since $\Gamma_{\min}(X)\leq 1$ and $\Gamma_{\max}(X)\geq 1$, we get the limit as
+$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}\big{(}\Gamma_{\max}(X),\Gamma_{\min}(X)\big{)}\bigg{)}$ (36)
+$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}(X)=\log\Gamma_{\max}(X).$ (37)
+The bound in [6, Theorem 35], [25, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
+Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergence of order $\alpha>1$.
+Theorem 8 (Pinsker’s inequality for $f_{\alpha}$-divergences).
 For two mutually absolutely continuous probability distributions $P_{X}$ and $Q_{X}$ and any $\alpha>1$, we have
@@
 where
-$g_{\alpha}(t)=\begin{cases}e^{2(\alpha-1)t^{2}}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha<2,\\ (4t^{2}+1)^{\alpha-1}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha\geq 2,\\ (1-t)^{1-\alpha}-1,&\text{ otherwise }\end{cases}$ (31)
+[…]
 Proof:
-The proof follows from bounding the joint range [2] of the two divergences and is presented in Appendix B. ∎
+The proof follows from bounding the joint range [1] of the divergences and is presented in Appendix A. ∎
 Remark 9.
@@
 As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$.
+Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
+$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=0.$ (40)
+In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$.
+Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by $g_{\alpha}(t)=\psi_{\alpha}(p^{*},t)$.
+However, it remains unclear if $p^{*}$ can be expressed in closed form.
+Due to this complication, the result above utilizes more rudimentary bounds in this case.
+Further, we invoke standard Pinsker’s inequality for relative entropy to derive the first case of (39).
+This bound is not the best we can do for this case.
+More specifically, the optimal Pinsker’s inequality has been found in [26] in implicit form.
+We rely on the more rudimentary standard Pinsker bound here for simplicity.
 Remark 10.
@@
 where $h_{2}(\alpha)\coloneqq(1+4\alpha^{-2})^{\alpha-1}-1$ and $h_{1}(\alpha)\coloneqq 2-\nicefrac{{2}}{{\alpha}}$.
-An example of the the presented bound on the joint range for $\alpha=4$ is shown in Figure 1.
+An example of the presented bound on the joint range for $\alpha=4$ is shown in Figure 1.
 Other values of $\alpha$ yield a visually similar result, with the discontinuity at $\nicefrac{{1}}{{\alpha}}$ moving closer to the origin as the value of $\alpha$ increases.
@@
 […]
-$\Gamma_{\max}=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (33)
-$\Gamma_{\min}=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (34)
+$\Gamma_{\max}(Y)=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (42)
+$\Gamma_{\min}(Y)=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (43)
 […]
 Corollary 11.
-For a discrete channel $P_{Y|X}$, a set $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ and some $\alpha>1$, let $\eta\coloneqq\eta_{TV}(P_{Y|X})$. We have
-$D_{f_{\alpha}}(P_{Y}||Q_{Y})\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(D_{f_{\alpha}}(P_{X}||Q_{X})).$ (35)
-Proof:
-Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$ and the inequality in Theorem 8. ∎
-This result stresses the relation between $\Gamma_{\max}$ and $\Gamma_{\min}$ as formulated in (33) and (34) and the contraction behavior of a channel: If $\Gamma_{\max}-\Gamma_{\min}$ is small, the channel significantly contracts the $f_{\alpha}$-divergence.
+For a discrete channel $P_{Y|X}$, distributions $P_{X},Q_{X}\in\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$, $P_{Y}=P_{Y|X}\circ P_{X}$, $Q_{Y}=P_{Y|X}\circ Q_{X}$ and some $\alpha>1$, we have
+$D_{f_{\alpha}}$ $(P_{Y}||Q_{Y})$ (44)
+$\leq\eta_{TV}(P_{Y|X},\mathcal{P})R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))$ (45)
+$\qquad\qquad\qquad\qquad\qquad\times g_{\alpha}^{-1}\Big{(}D_{f_{\alpha}}(P_{X}||Q_{X})\Big{)}.$ (46)
+Proof.
+Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$.
+∎
+This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence.
 If the difference is large, contraction will be small.
-What remains is to bound $\Gamma_{\max}$ and $\Gamma_{\min}$ such that an operational bound is obtained.
+What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ such that an operational bound on contraction is obtained.
 Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors.
-We present one such use-case specific bound in the following.
-VI Application: Post-Processing for RLDP
-In this section, we show how the above results can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [14, 25].
-Consider the Markov chain $W-X-Y$, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, i.e.,
-$\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (36)
-Further, assume that the transition kernel $P_{Y|X}$ does not satisfy any finite RLDP guarantee, that is, the channel matrix contains zero elements.
-From the post-processing property of RLDP, it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$.
-However, it is reasonable to assume that there exist channels $P_{Y|X}$ containing zero elements that will result in an improved guarantee of the composed system $P_{Y|W}$ if the output of the RLDP $P_{X|W}$ is applied to their input.
-In fact, using the results derived above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
+One such bound appears in the application of these results to privacy amplification, which we will present next.
+VI Privacy Amplification Bounds for RLDP
+In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [27, 5].
+Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [27] was recently suggested by Gilani et al.
+[5], where the authors refer to it as local Rényi differential privacy (LRDP).
+Note that the class of discrete mechanisms satisfying any (finite) RLDP guarantee for any $\alpha$ is equivalent to the class of discrete mechanisms satisfying some (finite) LDP guarantee.
+Specifically, this class only includes mechanisms with strictly positive entries in the mechanism matrix.
+i.e.,
+$\text{RLDP}^{(\alpha)}$ $(P_{X|W})$ (47)
+$\coloneqq\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (48)
+Further, assume that the channel $P_{Y|X}$ does not satisfy any finite RLDP guarantee.
+From the post-processing property of RLDP (which is just the DPI for Rényi-divergence), it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$.
+However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP channel $P_{X|W}$.
+Using our results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
 $D_{f_{\alpha}}$ $(P_{Y|W=w}||P_{Y|W=w^{\prime}})$ (49)
 $=D_{f_{\alpha}}(P_{Y|X}\circ P_{X|W=w}||P_{Y|X}\circ P_{X|W=w^{\prime}})$ (50)
-$\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}\biggl{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\biggr{)},$ (39)
-where $\eta\coloneqq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})$ is again a shorthand for the contraction coefficient of the total variation distance.
-Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can therefore bound the RLDP guarantee of the entire system by applying (4) as
-$\varepsilon^{*}\coloneqq\text{RLDP}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (40)
-where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence,
-$\varphi(\varepsilon_{f},\alpha)\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (41)
-Further, in this setup, the set of input distributions $\mathcal{P}$ is just the set of rows of the mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$.
-Hence, the quantities $\Gamma_{\max}$ and $\Gamma_{\min}$ reduce to
-$\Gamma_{\max}=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (42)
-$\Gamma_{\min}=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (43)
-While these quantities can be computed directly if both channels are perfectly known, it is useful to obtain a bound that only depends on the privacy guarantees of the channels.
-The following proposition provides such a bound, where the privacy of the second, non-RLDP channel is quantified by its pointwise maximal leakage (PML) [21].
-The proposition shows one way in which the increased flexibility of leakage assessment with PML can complement (L)DP guarantees.
-Proposition 12.
-If $P_{X|W}$ is a $N\times N$ randomized response mechanisms satisfying $\varepsilon$-LDP according to [31], then
-$\Gamma_{\max}\leq\biggl{(}1-\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (44)
-$\Gamma_{\min}\geq\biggl{(}1+\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (45)
-where $\ell_{w}(X\to y)$ denotes the PML from $X$ to an outcome $Y=y$ under the input distribution induced by $W=w$, i.e.,
-$\ell_{w}(X\to y)\coloneqq\log\frac{\max_{x\in\mathcal{X}}P_{Y|X=x}(y)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}.$ (46)
-Further, equality is achieved in both bounds if $P_{Y|X}$ has a zero entry in each column, i.e., if $\min_{x}P_{Y|X=x}(y)=0$ $\forall y\in\mathcal{Y}$.
-Proof:
-See Appendix C. ∎
-The following example illustrates this bound.
-Figure 2: RLDP and corresponding bound in Example 13.
+$\qquad\leq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))$ (51)
+$\qquad\qquad\qquad\quad\times g_{\alpha}^{-1}\Big{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\Big{)}.$ (52)
+Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can apply (4) to bound the RLDP guarantee of the entire system as
+$\varepsilon^{*}\coloneqq\text{RLDP}^{(\alpha)}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (53)
+where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence, and we define $\varphi:[0,\infty)\times(1,\infty)\to\mathbb{R}$ as
+$\varphi$ $(\varepsilon_{f},\alpha)$ (54)
+$\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max}(Y),\Gamma_{\min}(Y))g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (55)
+Further, in this setup, the set of input distributions $\mathcal{P}$ to the channel $P_{Y|X}$ is just the set of rows of the RLDP mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$.
+Hence, the quantities $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ reduce to
+$\Gamma_{\max}(Y)=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (56)
+$\Gamma_{\min}(Y)=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (57)
+The following example illustrates the above results.
+(a) $N=5$, $\varepsilon=\log 2$, $P_{Y|X}$
+(b) $N=5$, $\varepsilon=\log 6$, $P_{Y|X}$
+(c) $N=20$, $\varepsilon=\log 10$, $P_{Y|X}$
+(d) $N=100$, $\varepsilon=\log 10$, $Q_{Y|X}$
+Figure 2: RLDP guarantees and the corresponding bounds in Example 12.
 Example 12.
-Let $P_{X|W}$ be a $5\times 5$ randomized response mechanism in [31] satisfying $\log 6$-LDP, and $P_{Y|X}$ given by
-$P_{Y|X}=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0&0\\ 0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0\\ 0&0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ \nicefrac{{1}}{{2}}&0&0&0&\nicefrac{{1}}{{2}}\end{bmatrix}.$ (47)
-[…]
-Appendix A Proof of Proposition 3
-For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have
-$\eta_{\alpha}(P_{Y|X})=$ (48)
-$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (49)
-and the statement follows by constructing $P_{X}$ and $Q_{X}$ with full support such that in the limiting case to the support set boundaries, we get
-$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (50)
-Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, constructing this distribution proves $\eta_{\alpha}(P_{Y|X})=1$.
-Construction of $P_{X}$ and $Q_{X}$: For ease of notation, let $\mathcal{X}\coloneqq[N]$. Further, define the indicator function by
-$\delta_{i}(x)=\begin{cases}1,&\text{if }x=i\\ 0,&\text{otherwise}.\end{cases}$ (51)
-For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
-$P_{X}^{\gamma}(x)=(1-\gamma)\delta_{i}(x)+\frac{\gamma}{N-1}(1-\delta_{i}(x)),$ (52)
-and let
-$Q_{X}^{\gamma}(x)=\frac{1-\gamma}{2}(\delta_{i}(x)+\delta_{j}(x))+\frac{\gamma}{N-2}(1-\delta_{i}(x)-\delta_{j}(x)),$ (53)
-where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Then we get
-$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ $=\frac{\log\biggl{\{}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (54)
-$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (55)
-$=\frac{(1-\alpha)\log Q^{0}_{X}(i)}{(1-\alpha)\log Q^{0}_{X}(i)}=1,$ (56)
-where the second to last equality follows from the row-stochasticity of $P_{Y|X}$.
-By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$.
-Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$.
-Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof.
-∎
-Appendix B Proof of Theorem 8
+Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [28] satisfying $\varepsilon$-LDP, that is, let
+$P^{(N)}_{X|W=w}(x)=\begin{cases}\frac{e^{\varepsilon}}{N+e^{\varepsilon}-1},&\text{ if }x=w\\ \frac{1}{N+e^{\varepsilon}-1},&\text{ otherwise}.\end{cases}$ (58)
+We examine the privacy amplification of the channels $P_{Y|X}$ and $Q_{Y|X}$ defined for some $N$ and $i,j\in[N]$ by
+$P^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\nicefrac{{1}}{{2}},&\text{ if }i=j,\\ \nicefrac{{1}}{{2}},&\text{ if }i=j+1\mod N,\\ 0,&\text{otherwise}\end{cases}$ (59)
+and for even values of $N$,
+$Q^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\frac{2}{N},&\text{if }1\leq i,j\leq\frac{N}{2}\\ \frac{2}{N},&\text{if }\frac{N}{2}<i,j\leq N\\ 0,&\text{otherwise}\end{cases}$ (60)
+In this setup, the successive application of both mechanisms to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the complete system compared to $P^{(N)}_{X|W}$ alone.
+We can numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
+The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the composition $P^{(N)}_{X|W}\circ P^{(N)}_{Y|X}$ or $P^{(N)}_{Y|X}\circ P_{X|W}$.
+As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight.
+Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $P^{(N)}_{Y|X}$,
+$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{e^{\varepsilon}+1}{2}.$ (61)
+On the other hand, given $Q^{(N)}_{Y|X}$, we have
+$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{N+2e^{\varepsilon}-2}{N}.$ (62)
+That is, the contraction behavior of the channels $P^{(N)}_{Y|X}$—and hence its privacy amplification—is independent of the alphabet size $N$.
+On the other hand, the amplification of the channels with $Q^{(N)}_{Y|X}$ crucially depends on the alphabet size.
+Interestingly, as long as $\varepsilon<\infty$, we see that asymptotically, $\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}(Y)=1$, that is, the privacy amplification of the channels $Q^{(N)}_{Y|X}$ becomes independent of the privacy guarantee of the system $P^{(N)}_{X|W}$.
+In particular, post-processing a randomized response mechanism with any finite LDP guarantee with a channel of a structure similar to $Q^{(N)}_{Y|X}$ will result in a system with perfect privacy as $N\to\infty$.
+Example 12 shows that even sparse channel matrices can lead to significant privacy amplification if applied as a post-processing step to an LDP system.
+Moreover, as long as the number of non-zero elements in the channel matrix grows linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows.
+While these insights are so far limited to randomized response mechanisms, they are interesting for two main reasons:
+First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [29, 30].
+In this context, the results presented in this paper are encouraging.
+Specifically, if some form of LDP processing is done to the data, even a learning system without a formal LDP guarantee on its own might be able to significantly increase the privacy guarantee with respect to the data used to train it.
+Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [31]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.
+Future work in this direction should focus on more advanced tools and methods for bounding the values $\Gamma_{\max}$ and $\Gamma_{\min}$ for various (practical) systems.
+Obtaining such bounds paves the way for easy plug-and-play solutions for improving privacy guarantees of weak LDP systems at a minimal loss of utility.
+Appendix A Proof of Theorem 8
 From [6] and the fact that $D_{f_{\alpha}}$ is a monotone transform of the Rényi-divergence of the same order, we know that
@@
 $d_{f_{\alpha}}(p||q)\coloneqq\frac{p^{\alpha}}{q^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-q)^{\alpha-1}}-1.$ (64)
-From the definition of the total variation, we get
-$\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\biggl{\{}\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}\biggr{\}}-1.$ (59)
 Define
 $\psi_{\alpha}(p,t)\coloneqq\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}-1.$ (65)
-Taking derivatives with respect to $p$ yields
+From the definition of the total variation, we get $\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\{\psi_{\alpha}(p,t)\}$. Taking derivatives with respect to $p$ yields
 $\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}$ (66)
@@
 Together with the convexity of $\psi_{\alpha}$ in $p$, this implies that the maximum value of $\psi_{\alpha}$ for all $t\geq\nicefrac{{1}}{{\alpha}}$ will be achieved at $p=0$.
-At this point, we have
-$\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}.$ (64)
+At this point, we have $\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}$.
 Secondly, for $t<\nicefrac{{1}}{{\alpha}}$, we first consider $\alpha\geq 2$.
@@
 Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [4], so is $D_{f_{\alpha}}$.
-Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$)
-$\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$ (65)
-will continue to hold for $\alpha>1$. Utilizing (4), this yields
+Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$), $\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$, will continue to hold for $\alpha>1$.
+Utilizing (4), this yields
 $D_{f_{\alpha}}(P||Q)\geq\exp\big{(}2(\alpha-1)\text{TV}(P||Q)^{2}\big{)}-1.$ (69)
 Putting the derived bounds together to form $g_{\alpha}(t)$ finishes the proof. ∎
-Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
-$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}=0.$ (67)
-In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$.
-Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by
-$g_{\alpha}(t)=\psi_{\alpha}(p^{*},t).$ (68)
-However, it remains unclear if $p^{*}$ can be expressed in closed form. Due to this complication, the result above utilizes more rudimentary bounds in this case.
-Appendix C Proof of Proposition 12
-Note that for a randomized response mechanism according to [31], we can write for any two rows $w$, $w^{\prime}$,
-$P_{X|W=w}(x)=P_{X|W=w^{\prime}}(x)+\kappa(x),$ (69)
-where
-$\kappa(x)\in\{0,\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1},\frac{1-e^{\varepsilon}}{|\mathcal{X}|+e^{\varepsilon}-1}\}$ (70)
-and each non-zero element is taken exactly once. Therefore,
-$||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}.$ (71)
-Using this observation, we can write
-$\Gamma_{\max}$ $=\max_{w,w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w}(x)}$ (72)
-$=\max_{\kappa:\,||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}}\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)+\sum_{x}P_{Y|X=x}(y)\kappa(x)}$ (73)
-$=\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\biggl{(}\max_{x}P_{Y|X=x}(y)-\min_{x}P_{Y|X=x}(y)\biggr{)}}$ (74)
-$\leq\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\max_{x}P_{Y|X=x}(y)}.$ (75)
-Now, fix $y$ and let $w^{*}\in\mathcal{W}$ be the solution to the maximization over $\mathcal{W}$. By letting
-$\ell_{w^{*}}(X\to y)=\log\frac{\max_{x}P_{Y|X=x}(y)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{*}}(x)},$ (76)
-we can further simplify the above bound as
-$\Gamma_{\max}\leq\max_{y}\frac{1}{1-\frac{||\kappa||_{1}}{2}\exp\ell_{w^{*}}(X\to y)}=\biggl{(}1-\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1}.$ (77)
-The proof for $\Gamma_{\min}$ follows analogously. ∎
+Appendix B Proof of Proposition 3
+For some fixed $P_{Y|X}$, let $P_{Y}=P_{Y|X}\circ P_{X}$ and $Q_{Y}=P_{Y|X}\circ Q_{X}$. For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have
+$\eta_{\alpha}(P_{Y|X})=$ (70)
+$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\bigg{\{}\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (71)
+and the statement follows by constructing $P_{X}$ and $Q_{X}$ with full support such that in the limiting case to the support set boundaries, we get
+$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (72)
+Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this then implies $\eta_{\alpha}(P_{Y|X})=1$.
+Construction of $P_{X}$ and $Q_{X}$: For ease of notation, let $\mathcal{X}\coloneqq[N]$. Further, define the indicator function by
+$\delta_{i}(x)=\begin{cases}1,&\text{if }x=i\\ 0,&\text{otherwise}.\end{cases}$ (73)
+For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
+$P_{X}^{\gamma}(x)=(1-\gamma)\delta_{i}(x)+\frac{\gamma}{N-1}(1-\delta_{i}(x)),$ (74)
+and let
+$Q_{X}^{\gamma}(x)=\frac{1-\gamma}{2}(\delta_{i}(x)+\delta_{j}(x))+\frac{\gamma}{N-2}(1-\delta_{i}(x)-\delta_{j}(x)),$ (75)
+where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Let $\mathcal{S}_{i}\coloneqq\textnormal{supp}(P_{Y|X=i})$. Then we get
+$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ (76)
+$=\frac{\log\biggl{\{}\sum_{y\in\mathcal{S}_{i}}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (77)
+$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\mathcal{S}_{i}}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (78)
+$=\frac{(1-\alpha)\log Q^{0}_{X}(i)}{(1-\alpha)\log Q^{0}_{X}(i)}=1,$ (79)
+where the second to last equality follows from the row-stochasticity of $P_{Y|X}$.
+By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$.
+Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$.
+Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof.
+∎

### v4

@@
 We examine the privacy amplification of channels that do not necessarily satisfy any LDP guarantee by analyzing their contraction behavior in terms of $f_{\alpha}$-divergence, an $f$-divergence related to Rényi-divergence via a monotonic transformation.
-We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by Harremoës and Vajda [1].
+We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by 5773031.
 The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$.
@@
 We take steps to answer this question in this paper.
-In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [2, Chapter 33]).
-For pure LDP, the divergence measure of interest is the Rényi-divergence [3, 4] of order $\infty$, since
+In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [Polyanskiy_Wu_2025, Chapter 33]).
+For pure LDP, the divergence measure of interest is the Rényi-divergence [renyi1961entropy, van2014renyi] of order $\infty$, since
 $\max_{y}\log\frac{P_{X|W=w}(y)}{P_{X|W=w^{\prime}}(y)}=\lim_{\alpha\to\infty}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}}).$ (1)
-Gilani et al.
-[5] suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$.
+gilani2024unifying suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$.
 With this framework at hand, studying the privacy amplification of a channel becomes equivalent to studying SDPIs for Rényi-divergences.
@@
 Our key contributions are:
-  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on privacy amplification in Section III.
+  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on the contraction of a channel in Section III.
   • We develop new tools for bounding the $f_{\alpha}$-divergence in terms of total variation distance in Section IV. Specifically, we provide:
     • an improved and often tight version of Pinsker’s inequality for $f_{\alpha}$-divergence (for $\alpha>1$);
-    • a simplification of the recently presented proof of Binette’s inequality [7] in [8] based on $E_{\gamma}$-divergence.
+    • a simplification of the recently presented proof of Binette’s inequality [8630660] in [Hirche:2023caq] based on $E_{\gamma}$-divergence.
   • Using these tools, we analyze the contraction of $f_{\alpha}$-divergence and show how input distribution constraints can yield meaningful bounds in Section V.
@@
 I-B Related Work
-The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [2, 9, 10, 11, 12, 13].
-For Rényi-divergences, Polyanskiy and Wu [14] provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$.
-Recently, Jin et al.
-[15] show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid.
-In the context of privacy, Zamanlooy and Asoodeh [11] show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee.
-These bounds are further applied to mixing times in [16].
+The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [Polyanskiy_Wu_2025, raginsky2016strong, 10.1134/S0032946020020015, 10206578, anantharam2013maximal, ahlswede1976spreading].
+For Rényi-divergences, polyanskiy2015dissipation provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$.
+Recently, 10619367 show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid.
+In the context of privacy, 10206578 show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee.
+These bounds are further applied to mixing times in [zamanlooy2024mathrm].
 Related, it is shown in [9517999] that the LDP guarantee of a mechanism is closely related to its $f$-divergence contraction properties.
-The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [6] and subsequent works, including an applications to SDPIs in [18].
+The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [7552457] and subsequent works, including an applications to SDPIs in [sason2019data].
+The privacy amplification of various mechanisms in the central RDP setting is studied in the line of work [balle2018privacy, balle2019privacy, balle2020privacy].
 II Preliminaries
@@
 II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
-For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [19] as
+For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [csiszar1967information] as
 $D_{f}(P||Q)=\mathbb{E}_{Q}\bigg[f\bigg(\frac{dP}{dQ}\bigg)\bigg].$ (2)
-The Rényi-divergence [3] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [2]:
+The Rényi-divergence [renyi1961entropy] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [Polyanskiy_Wu_2025]:
 Definition 1 ($f_{\alpha}$-divergence).
@@
 $D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\biggl\{D_{f_{\alpha}}(P||Q)+1\biggr\}.$ (4)
-We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [14, Section 3.2].
+We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [polyanskiy2015dissipation, Section 3.2].
 Definition 2 ($E_{\gamma}$-divergence, see, e.g., [7552457]).
@@
 $E_{\gamma}(P||Q)=\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-\gamma Q(x)|-\frac{1}{2}|1-\gamma|.$ (5)
-For twice differentiable $f$, the relation shown in [20],
+For twice differentiable $f$, the relation shown in [cohen1998comparisons],
 $D_{f}(P||Q)=\int_{0}^{\infty}E_{\gamma}(P||Q)f^{\prime\prime}(\gamma)d\gamma,$ (6)
@@
 $\mathbb{E}\big[f(X)\big]=\mathbb{E}\big[g(Y)\big]=0,\quad\mathbb{E}\big[f^{2}(X)\big]=\mathbb{E}\big[g^{2}(Y)\big]=1.$ (13)
-The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, in particular, it is shown in, e.g., [2, Theorem 33.6(c)] that
+The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, and it is shown in, e.g., [Polyanskiy_Wu_2025, Thm. 33.6(c)] that
 $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\rho_{m}^{2}(X;Y)\eqqcolon s(P_{Y|X}).$ (14)
-It was shown by Ahlswede and Gács [13] that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$.
+It was shown by ahlswede1976spreading that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$.
 Together with the data processing inequality for $f$-divergences, this implies that whenever (11) holds, we have $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
-Next, it was shown in [15, Corollary 2] that
+Next, it was shown in [10619367, Corollary 2] that
 $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\geq\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}}).$ (15)
@@
 Proposition 3 states that a channel does not contract the Rényi-divergence whenever there exist two input symbols that cause conditional distributions with disjoint support at the channel output.
-This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [21]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
+This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [diestel2017]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
 $W=\{(x_{i},x_{j})\in\mathcal{X}^{2}\mid\exists\,y\in\mathcal{Y}:x_{i}\in\mathcal{X}^{(y)}\wedge x_{j}\in\mathcal{X}^{(y)}\},$ (16)
@@
 That is, a missing edge (compared to the complete version) indicates the existence of a pair of symbols that can be sent over the channel without being confused with each other.
-The confusion graph–along with its graph-theoretic dual–is extensively used in zero-error information theory, see e.g.
-[22, 23] for a detailed account of results and methods.
+The confusion graph—along with its graph-theoretic dual—is extensively used in zero-error information theory, see e.g., [1056798, 720537] for a detailed account of results and methods.
 Note that a fully connected confusion graph is not necessarily equivalent to a strictly positive channel matrix.
@@
 As we will see in Section VI, this can be useful in cases in which partial knowledge about the input distributions to a channel is available.
-[…]
+In this section, we present a generalization of the reverse Pinker’s inequality and Pinsker’s inequality for Rényi-divergences.
+These inequalities will then be used to derive “SDPI-style” bounds in Section V.
+In most of what follows, we will work with the $f_{\alpha}$-divergence as a substitute for Rényi-divergence.
+Due to the monotonic transformation in (4), conversion between the two can be done easily at any point.
 The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions.
-It was first shown by Binette [7] for general $f$-divergences.
+It was first shown by 8630660 for general $f$-divergences.
 Here, we present a proof that utilizes the integral representation of $f$-divergences via the $E_{\gamma}$-divergence in (6).
-We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche and Tomamichel [8, Proposition 5.2].
+We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche:2023caq.
 However, the proof presented below slightly simplifies arguments, and we include it here as we believe it to be instructive.
-Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [7, Theorem 1]).
+Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [8630660, Theorem 1]).
 Let $\alpha>1$. For any two distributions $P_{X},Q_{X}\in\mathcal{P}$, define
@@
 What remains is to bound $E_{\gamma}$ on the interval $I_{\geq 0}$, which yields an upper bound on $D_{f_{\alpha}}$ by (6).
-To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [14].
+To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [polyanskiy2015dissipation].
 Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$.
@@
 $=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}(X)=\log\Gamma_{\max}(X).$ (37)
-The bound in [6, Theorem 35], [25, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
+The bound in [7552457, Theorem 35], [7360766, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
 Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergence of order $\alpha>1$.
 Theorem 8 (Pinsker’s inequality for $f_{\alpha}$-divergences).
-For two mutually absolutely continuous probability distributions $P_{X}$ and $Q_{X}$ and any $\alpha>1$, we have
-$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl{(}\text{TV}(P_{X}||Q_{X})\biggr{)}$ (38)
+For any $\alpha>1$ and any two mutually absolutely continuous probability measures $P_{X}\ll\gg Q_{X}$, we have
+$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl(\text{TV}(P_{X}||Q_{X})\biggr),$ (38)
 where
@@
 Remark 9.
-For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$.
+For any $\alpha>1$, the bound in Theorem 8 is attained by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$.
 Hence, it is tight for this range of values.
@@
 In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g.
-presented in [2].
-As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$.
+presented in [Polyanskiy_Wu_2025].
+As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ holds for all $t>0$.
 Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
@@
 V Strong Data Processing with Pinsker
-[…]
+The above results can now be used to reason about the data processing properties of channels by considering the presented inequalities “across” a given channel.
+That is, we can modify the bound presented in Theorem 6 to yield an upper bound on $D_{f_{\alpha}}(P_{Y}||Q_{Y})$ in terms of $\text{TV}(P_{X}||Q_{X})$.
+To this end, consider
 $\Gamma_{\max}(Y)=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (42)
 $\Gamma_{\min}(Y)=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)},$ (43)
-[…]
+as the “cross-channel” modification of the crucial extremal quantities in Theorem 6. We get the following result.
 Corollary 11.
@@
 Proof.
-Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$.
+Follows from Theorem 6 with $P_{Y}=P_{Y|X}\circ P_{X}$ and $Q_{Y}=P_{Y|X}\circ Q_{X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$.
 ∎
-This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence.
+This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel in terms of Rényi-divergence: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence.
 If the difference is large, contraction will be small.
-What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ such that an operational bound on contraction is obtained.
+What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ so that an operational bound on this contraction is obtained.
 Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors.
-One such bound appears in the application of these results to privacy amplification, which we will present next.
+One such bound can be obtained in the application of the above results to privacy amplification, which we will present in the following section.
 VI Privacy Amplification Bounds for RLDP
-In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [27, 5].
-Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [27] was recently suggested by Gilani et al.
-[5], where the authors refer to it as local Rényi differential privacy (LRDP).
+In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [mironov2017renyi, gilani2024unifying].
+Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [mironov2017renyi] was recently suggested by gilani2024unifying, where the authors refer to it as local Rényi differential privacy (LRDP).
 Note that the class of discrete mechanisms satisfying any (finite) RLDP guarantee for any $\alpha$ is equivalent to the class of discrete mechanisms satisfying some (finite) LDP guarantee.
@@
 From the post-processing property of RLDP (which is just the DPI for Rényi-divergence), it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$.
-However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP channel $P_{X|W}$.
-Using our results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
+However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP mechanism $P_{X|W}$.
+Using the results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
 $D_{f_{\alpha}}$ $(P_{Y|W=w}||P_{Y|W=w^{\prime}})$ (49)
@@
 The following example illustrates the above results.
-(a) $N=5$, $\varepsilon=\log 2$, $P_{Y|X}$
-(b) $N=5$, $\varepsilon=\log 6$, $P_{Y|X}$
-(c) $N=20$, $\varepsilon=\log 10$, $P_{Y|X}$
-(d) $N=100$, $\varepsilon=\log 10$, $Q_{Y|X}$
+(a) $N=5$, $\varepsilon=\log 2$, $k=2$
+(b) $N=5$, $\varepsilon=\log 6$, $k=2$
+(c) $N=20$, $\varepsilon=\log 10$, $k=2$
+(d) $N=100$, $\varepsilon=\log 10$, $k=50$
 Figure 2: RLDP guarantees and the corresponding bounds in Example 12.
 Example 12.
-Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [28] satisfying $\varepsilon$-LDP, that is, let
+Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [extremalmechanismLong] satisfying $\varepsilon$-LDP, that is, let
 $P^{(N)}_{X|W=w}(x)=\begin{cases}\frac{e^{\varepsilon}}{N+e^{\varepsilon}-1},&\text{ if }x=w\\ \frac{1}{N+e^{\varepsilon}-1},&\text{ otherwise}.\end{cases}$ (58)
-We examine the privacy amplification of the channels $P_{Y|X}$ and $Q_{Y|X}$ defined for some $N$ and $i,j\in[N]$ by
-$P^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\nicefrac{{1}}{{2}},&\text{ if }i=j,\\ \nicefrac{{1}}{{2}},&\text{ if }i=j+1\mod N,\\ 0,&\text{otherwise}\end{cases}$ (59)
-and for even values of $N$,
-$Q^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\frac{2}{N},&\text{if }1\leq i,j\leq\frac{N}{2}\\ \frac{2}{N},&\text{if }\frac{N}{2}<i,j\leq N\\ 0,&\text{otherwise}\end{cases}$ (60)
-In this setup, the successive application of both mechanisms to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the complete system compared to $P^{(N)}_{X|W}$ alone.
-We can numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
-The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the composition $P^{(N)}_{X|W}\circ P^{(N)}_{Y|X}$ or $P^{(N)}_{Y|X}\circ P_{X|W}$.
-As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight.
-Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $P^{(N)}_{Y|X}$,
-$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{e^{\varepsilon}+1}{2}.$ (61)
-On the other hand, given $Q^{(N)}_{Y|X}$, we have
-$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{N+2e^{\varepsilon}-2}{N}.$ (62)
-That is, the contraction behavior of the channels $P^{(N)}_{Y|X}$—and hence its privacy amplification—is independent of the alphabet size $N$.
-On the other hand, the amplification of the channels with $Q^{(N)}_{Y|X}$ crucially depends on the alphabet size.
-Interestingly, as long as $\varepsilon<\infty$, we see that asymptotically, $\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}(Y)=1$, that is, the privacy amplification of the channels $Q^{(N)}_{Y|X}$ becomes independent of the privacy guarantee of the system $P^{(N)}_{X|W}$.
-In particular, post-processing a randomized response mechanism with any finite LDP guarantee with a channel of a structure similar to $Q^{(N)}_{Y|X}$ will result in a system with perfect privacy as $N\to\infty$.
+For some $m,k>0$, let $N=mk$. We examine the privacy amplification of the channels $Q^{(m,k)}_{Y|X}$ defined by
+$Q_{Y|X}^{(m,k)}\coloneqq\text{diag}(\underbrace{U^{(k)},\dots,U^{(k)}}_{m\text{ times}}),$ (59)
+where $U^{(k)}=\frac{1}{k}\mathbbm{1}_{k}\mathbbm{1}_{k}^{T}$ is a “uniform” block of size $k\times k$.
+In this setup, the successive application of $P_{X|W}^{(N)}$ and $Q_{Y|X}^{(m,k)}$ to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the composed system compared to $P^{(N)}_{X|W}$ alone.
+Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $Q^{(m,k)}_{Y|X}$, we have
+$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=1+\frac{e^{\varepsilon}-1}{k}=1+\frac{m(e^{\varepsilon}-1)}{N}.$ (60)
+Interestingly, we observe that for any $\varepsilon>0$, and any fixed ratio of non-zero elements $\nicefrac{{k}}{{N}}=\nicefrac{{1}}{{m}}$, we have
+$\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}=1,$ (61)
+which implies that for any $\alpha>1$, we have
+$\varepsilon^{*}=\text{RLDP}^{(\alpha)}\Big(Q^{(\nicefrac{{N}}{{k}},k)}_{Y|X}\circ P_{X|W}^{(N)}\Big)\longrightarrow 0.$ (62)
+That is, post-processing a randomized response mechanism of any finite LDP guarantee with a channel constructed in the above way will result in a system with perfect privacy as $N\to\infty$ whenever the ratio of non-zero elements remains fixed as $N$ grows.
+We can further numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
+The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the compositions for different choices of $N$ and $k$.
+As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight in the parameter $\alpha$.
 Example 12 shows that even sparse channel matrices can lead to significant privacy amplification if applied as a post-processing step to an LDP system.
-Moreover, as long as the number of non-zero elements in the channel matrix grows linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows.
+Moreover, as long as the number of non-zero elements in the channel matrix grows at least linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows.
 While these insights are so far limited to randomized response mechanisms, they are interesting for two main reasons:
-First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [29, 30].
+First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [10.1145/3433638, 10.5555/3454287.3455674].
 In this context, the results presented in this paper are encouraging.
 Specifically, if some form of LDP processing is done to the data, even a learning system without a formal LDP guarantee on its own might be able to significantly increase the privacy guarantee with respect to the data used to train it.
-Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [31]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.
+Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [dwork2019exposeyourepsilons]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.
 Future work in this direction should focus on more advanced tools and methods for bounding the values $\Gamma_{\max}$ and $\Gamma_{\min}$ for various (practical) systems.
@@
 $\geq(1+4t^{2})^{\alpha-1}-1.$
-Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [4], so is $D_{f_{\alpha}}$.
+Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [van2014renyi], so is $D_{f_{\alpha}}$.
 Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$), $\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$, will continue to hold for $\alpha>1$.
@@
 For some fixed $P_{Y|X}$, let $P_{Y}=P_{Y|X}\circ P_{X}$ and $Q_{Y}=P_{Y|X}\circ Q_{X}$. For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have
-$\eta_{\alpha}(P_{Y|X})=$ (70)
+$\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=$ (70)
 $\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\bigg\{\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr\}}{\log\biggl\{\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr\}}$ (71)
@@
 $\frac{\log\biggl\{\sum_{y\in\mathcal{Y}}P_{Y}(y)^{\alpha}Q_{Y}(y)^{1-\alpha}\biggr\}}{\log\biggl\{\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr\}}=1.$ (72)
-Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this then implies $\eta_{\alpha}(P_{Y|X})=1$.
+Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, this implies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
 Construction of $P_{X}$ and $Q_{X}$: For ease of notation, let $\mathcal{X}\coloneqq[N]$. Further, define the indicator function by
 $\delta_{i}(x)=\begin{cases}1,&\text{if }x=i\\ 0,&\text{otherwise}.\end{cases}$ (73)
-For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
+For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a uniformly perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
 $P_{X}^{\gamma}(x)=(1-\gamma)\delta_{i}(x)+\frac{\gamma}{N-1}(1-\delta_{i}(x)),$ (74)
-and let
+and similarly, let
 $Q_{X}^{\gamma}(x)=\frac{1-\gamma}{2}(\delta_{i}(x)+\delta_{j}(x))+\frac{\gamma}{N-2}(1-\delta_{i}(x)-\delta_{j}(x)),$ (75)